Determine the molar solubility of Fe(OH)2 in pure water. Ksp for Fe(OH)2)= 4.87 × 10-17.

Please explain your answer and I will rate 5 stars!

Thanks!
Answer #1

Fe(OH)2 (s) <=> Fe2+ + 2 OH-

let x = mol/L of Fe(OH)2 that dissolve : this will give x mol/L Fe2+ and 2 mol/L OH-


Ksp = 4.87 x 10^-17 = [Fe2+][OH-]^2 = (x)(2x)^2 = 4 x^3

x = molar solubility =2.30 x 10^-6 M
Answer #2

Fe(OH)2 (s) <-> Fe2+ + 2 OH- let x = mol/L of Fe(OH)2 that dissolve : this will give x mol/L Fe2+ and 2 molL OH- Ksp 4.87 x 1